# Exponential decay property for eigenfunctions of quantum walks

### v1

@@ Abstract
AMS 2020 Classification: 47B02,  47B15,  47B93,  81Q35.
Keywords: Eigenfunction, Eigenvalue, Exponential decay, Quantum walk, Unitary operator.
1 Introduction
@@
[…]
@@ 2 Set up and main result
  1. 1.
    It follows that $\sigma_{\mathrm{d}}(U)\neq\emptyset$.
  2. 2.
    A unitary operator $U$ finitely propagates with respect to a non-negative self-adjoint operator $A$. Moreover, $A$ is unbounded.
@@
Under Assunption 2.1, for any $\psi\in\mathrm{Ker}(U-\lambda)\setminus\{0\}$ with $\lambda\in\sigma_{\mathrm{d}}(U)$, $e^{\delta A}\psi\in\mathcal{H}$ for any $\delta>0$ such that $2\sinh(b\delta)<d(\lambda)$.
@@
A non-negativity in the second part of Assumption 2.1 is not essential. However, for simplicity, we assume the non-negativity of $A$ in this paper.
@@
$\|Uf-\lambda f\|\geq\{d(\lambda)-\epsilon\}\|f\|,\ d(\lambda):=\mathrm{dist}\{\lambda,\sigma_{\mathrm{ess}}(U)\},$
@@
$a\|g_{R}\|<\|Ug_{R}-\lambda g_{R}\|\leq\|Uf_{R}-\lambda f_{R}\|+\|(U-\lambda)E_{U}(X)f_{R}\|<d(\lambda)-\epsilon+2\|E_{U}(X)f_{R}\|.$
@@
For any $R>0$, we define the cut-off function $\chi_{R}$ as
$\chi_{R}(r):=\begin{cases}0,\quad r\in[0,R),\\ 1,\quad r\in[R,\infty).\end{cases}$
Then, $e^{A}[U,\chi_{R}(A)]$ is bounded on $\mathcal{H}$ and $\|e^{A}[U,\chi_{R}(A)]\|\leq e^{R+b}+e^{R}.$
@@
where if $R-b\leq 0,$ we set $E_{A}([0,R-b)=0$ and $E_{A}([R-b,R))=E_{A}([0,R)).$ Thus, for any $\psi\in\mathcal{H},$ it follows that $[U,\chi_{R}(A)]\psi\in D(e^{A})$ and
$\|e^{A}[U,\chi_{R}(A)]\psi\|\leq(e^{R+b}+e^{R})\|\psi\|.$
@@
$\Lambda(r):=\displaystyle\sum_{n=1}^{\infty}b\delta n\mathbb{I}_{B_{n}}(r),\quad\Lambda_{N}(r):=\begin{cases}\displaystyle\sum_{n=1}^{N}b\delta n\mathbb{I}_{B_{n}}(r),\ &r\in[0,Nb),\\ b\delta N,&r\in[Nb,\infty),\end{cases}$ (2.1)
where $B_{n}:=[(n-1)b,nb)\subset\mathbb{R}$ and $\mathbb{I}_{B_{n}}$ is the characteristic function of $B_{n}.$ Then, $\Lambda$ approximates a function $f(r):=\delta r,\ (r\in[0,\infty))$ from the above and $\Lambda_{N}$ is a cut-off function of $\Lambda.$ We note that $[U,e^{\Lambda_{N}(A)}]e^{-\Lambda_{N}(A)}$ is bounded for any $N\in\mathbb{N}.$
@@
$\|[U,e^{\Lambda_{N}(A)}]e^{-\Lambda_{N}(A)}\|\leq 2\sinh(b\delta).$
@@
[…]
$e^{t\Lambda_{N}(A)}[U,\Lambda_{N}(A)]e^{-t\Lambda_{N}(A)}$ $=e^{t\Lambda_{N}(A)}\{U\Lambda_{N}(A)-\Lambda_{N}(A)U\}E_{A}(B_{1})$
$+\displaystyle\sum_{m=2}^{N-1}e^{t\Lambda_{N}(A)}\{U\Lambda_{N}(A)-\Lambda_{N}(A)U\}e^{-t\Lambda_{N}(A)}E_{A}(B_{m})$
$+e^{t\Lambda_{N}(A)}\{U\Lambda_{N}(A)-\Lambda_{N}(A)U\}e^{-t\Lambda_{N}(A)}E_{A}(B_{N})$
@@
$+e^{t\Lambda_{N}(A)}\{U\Lambda_{N}(A)-\Lambda_{N}(A)U\}e^{-t\Lambda_{N}(A)}E_{A}([(N+1)b,\infty))$
$=:\mathrm{I}+\mathrm{II}+\mathrm{III}+\mathrm{IV}+\mathrm{V}.$
@@
$=E_{A}(B_{2})e^{2tb\delta}(Ub\delta-2b\delta U)e^{-tb\delta}E_{A}(B_{1})$
$=-\delta e^{tb\delta}E_{A}(B_{2})UE_{A}(B_{1}).$
@@
$\mathrm{II}$ $=\displaystyle\sum_{m=2}^{N-1}\{E_{A}(B_{m-1})+E_{A}(B_{m})+E_{A}(B_{m+1})\}e^{t\Lambda_{N}(A)}\{U\Lambda_{N}(A)-\Lambda_{N}(A)U\}e^{-\Lambda_{N}(A)}E_{A}(B_{m})$
$=\displaystyle\sum_{m=2}^{N-1}\Big{[}E_{A}(B_{m-1})e^{tb\delta(m-1)}\{Ub\delta m-b\delta(m-1)U\}e^{-tb\delta m}E_{A}(B_{m})$
$\quad+E_{A}(B_{m+1})e^{tb\delta(m+1)}\{Ub\delta m-b\delta(m+1)U\}e^{-tb\delta m}E_{A}(B_{m})\Big{]}$
$=\delta b\displaystyle\sum_{m=2}^{N-1}\Big{[}e^{-tb\delta}E_{A}(B_{m-1})UE_{A}(B_{m})-e^{tb\delta}E_{A}(B_{m+1})UE_{A}(B_{m})\Big{]}.$
@@
$\mathrm{III}$ =$=\{E_{A}(B_{N-1})+E_{A}(B_{N})+E_{A}(B_{N+1})\}e^{t\Lambda_{N}(A)}\{U\Lambda_{N}(A)-\Lambda_{N}(A)U\}e^{-t\Lambda_{N}(A)}E_{A}(B_{N})$
$=E_{A}(B_{N-1})e^{tb\delta(N-1)}\{Ub\delta N-b\delta(N-1)U\}e^{-tb\delta N}E_{A}(B_{N})$
$\quad+E_{A}(B_{N+1})e^{tb\delta(N+1)}\{Ub\delta N-b\delta(N+1)U\}e^{-tb\delta N}E_{A}(B_{N})$
$=b\delta e^{-tb\delta}E_{A}(B_{N-1})UE_{A}(B_{N})-b\delta e^{tb\delta}E_{A}(B_{N+1})UE_{A}(B_{N}).$
The forth term $\mathrm{IV}$ can be calculated as follows:
$\mathrm{IV}$ =$=\{E_{A}(B_{N})+E_{A}(B_{N+1})+E_{A}(B_{N+2})\}e^{t\Lambda_{N}(A)}\{U\Lambda_{N}(A)-\Lambda_{N}(A)U\}e^{-t\Lambda_{N}(A)}E_{A}(B_{N+1})$
$=E_{A}(B_{N})e^{tb\delta N}\{Ub\delta(N+1)-b\delta NU\}e^{-tb\delta(N+1)}E_{A}(B_{N+1})$
$=b\delta e^{-tb\delta}E_{A}(B_{N})UE_{A}(B_{N+1}).$
Lastly, the fifth term $\mathrm{V}$ can be calculated as follows:
$\mathrm{V}$ $=E_{A}([Nb,\infty))e^{t\Lambda_{N}(A)}\{U\Lambda_{N}(A)-\Lambda_{N}(A)U\}e^{-t\Lambda_{N}(A)}E_{A}([(N+1)b,\infty)$
$=E_{A}([Nb,\infty)e^{t\Lambda_{N}(A)}(Ub\delta N-b\delta NU)e^{-t\Lambda_{N}(A)}E_{A}([(N+1)b,\infty)$
@@
$[U,e^{\Lambda_{N}(A)}]e^{-\Lambda_{N}(A)}$ $=b\delta\displaystyle\int_{0}^{1}e^{-tb\delta}\mathrm{d}t\cdot\displaystyle\sum_{m=2}^{N+1}E_{A}(B_{m-1})UE_{A}(B_{m})-b\delta\displaystyle\int_{0}^{1}e^{tb\delta}\mathrm{d}t\cdot\displaystyle\sum_{m=1}^{N}E_{A}(B_{m+1})UE_{A}(B_{m})$
$=(1-e^{-b\delta})\displaystyle\sum_{m=2}^{N+1}E_{A}(B_{m-1})UE_{A}(B_{m})-(e^{b\delta}-1)\displaystyle\sum_{m=1}^{N}E_{A}(B_{m+1})UE_{A}(B_{m}).$
@@
$\|[U,e^{\Lambda_{N}(A)}]e^{-\Lambda_{N}(A)}\psi\|^{2}$ $=\|(1-e^{-b\delta})\displaystyle\sum_{m=2}^{N+1}E_{A}(B_{m-1})UE_{A}(B_{m})\psi-(e^{b\delta}-1)\displaystyle\sum_{m=1}^{N}E_{A}(B_{m+1})UE_{A}(B_{m})\psi\|^{2}$
$=(1-e^{-b\delta})^{2}\displaystyle\sum_{m=2}^{N+1}\|E_{A}(B_{m-1})UE_{A}(B_{m})\psi\|^{2}+(e^{b\delta}-1)^{2}\displaystyle\sum_{m=1}^{N}\|E_{A}(B_{m+1})UE_{A}(B_{m})\psi\|^{2}$
$\quad-2(1-e^{-b\delta})(e^{b\delta}-1)\displaystyle\sum_{m=2}^{N+1}\displaystyle\sum_{n=1}^{N}\mathrm{Re}\langle E_{A}(B_{m-1})UE_{A}(B_{m})\psi,E_{A}(B_{n+1})UE_{A}(B_{n})\psi\rangle$
$=(1-e^{-b\delta})^{2}\displaystyle\sum_{m=2}^{N+1}\|E_{A}(B_{m-1})UE_{A}(B_{m})\psi\|^{2}+(e^{b\delta}-1)^{2}\displaystyle\sum_{m=1}^{N}\|E_{A}(B_{m+1})UE_{A}(B_{m})\psi\|^{2}$
$\quad-2(1-e^{-b\delta})(e^{b\delta}-1)\displaystyle\sum_{n=2}^{N}\mathrm{Re}\langle E_{A}(B_{n})UE_{A}(B_{n+1})\psi,E_{A}(B_{n})UE_{A}(B_{n-1})\psi\rangle$
$\leq(1-e^{-b\delta})^{2}\displaystyle\sum_{m=2}^{N}\|E_{A}(B_{m})\psi\|^{2}+(e^{b\delta}-1)^{2}\displaystyle\sum_{m=1}^{N}\|E_{A}(B_{m})\psi\|^{2}$
$\quad+(1-e^{-b\delta})(e^{b\delta}-1)\displaystyle\sum_{n=2}^{N}\{\|E_{A}(B_{n})UE_{A}(B_{n+1})\psi\|^{2}+\|E_{A}(B_{n})UE_{A}(B_{n-1})\psi\|^{2}\}$
$\leq(1-e^{b\delta})^{2}\|\psi\|^{2}+(e^{b\delta}-1)^{2}\|\psi\|^{2}+2(1-e^{-b\delta})(e^{b\delta}-1)\|\psi\|^{2}$
=$=\{(1-e^{-b\delta})+(e^{b\delta}-1)\}^{2}\|\psi\|^{2}$
$=(e^{b\delta}-e^{-b\delta})^{2}\|\psi\|^{2}.$
@@
We choose $\epsilon>0$ as $\epsilon:=[d(\lambda)-2\sinh(b\delta)]/2.$ Then, by Lemma 2.1, there exists $R>0$ such that for any $f\in\mathcal{H}$ with $E([0,R))f=0,$ we have
@@
We take $\psi\in\mathrm{Ker}(U-\lambda)\setminus\{0\}$ with $\lambda\in\sigma_{\mathrm{d}}(U)$. For $R$ and $b$, there exists $N_{0}\in\mathbb{N}$ such that $R<N_{0}b$. Then we set $f_{N}:=e^{\Lambda_{N}(A)}\chi_{R}(A)\psi,\ (N\geq N_{0}).$ Since $E_{A}([0,R))f=0$, we have the following for arbitrary $N\geq N_{0}$:
@@
$Uf_{N}-\lambda f_{N}=[U,e^{\Lambda_{N}(A)}\chi(A)]\psi=[U,e^{\Lambda_{N}(A)}]\chi_{R}(A)\psi+e^{\Lambda_{N}(A)}[U,\chi_{R}(A)]\psi.$ (2.4)
@@
$\|e^{\Lambda_{N}(A)}[U,\chi_{R}(A)]\psi\|\leq(e^{R}+e^{R+b})\|\psi\|,$
[…]
$\|[U,e^{\Lambda_{N}(A)}]\chi_{R}(A)\psi\|=\|[U,e^{\Lambda_{N}(A)}]e^{-\Lambda_{N}(A)}e^{\Lambda_{N}(A)}\chi_{R}(A)\psi\|\leq 2\sinh(b\delta)\|f_{N}\|.$
@@
$\|Uf_{N}-\lambda f_{N}\|\leq(e^{R}+e^{R+b})\|\psi\|+2\sinh(b\delta)\|f_{N}\|.$
[…]
$\displaystyle\frac{d(\lambda)-2\sinh(b\delta)}{2}\|f_{N}\|\leq(e^{R}+e^{R+b})\|\psi\|.$ (2.5)
[…]
3 Application

### v2

@@ Abstract.
Key words and phrases:
Eigenfunction, Eigenvalue, Exponential decay, Quantum walk, Unitary operator
1991 Mathematics Subject Classification:
Primary 81Q35; Secondary 47B02, 47B15, 47B93
1. Introduction
@@
First, we establish (1.1) under a general setting in Section 2. Since we treat exponential operators of unbounded operators, we have to introduce suitable cut-off functions to avoid domain problems. For the proof, we mainly follow the methods presented by Yafaev [38] concerned the first-order differential systems involving Dirac operators. In our case, the derivative of functions are replaced by commutators. To analyze commutators is the crucial part.
In proofs, instead of $A$, we introduce another operator $\Lambda(A)$ which is step-like and approximates $A$ from above (see (2.1)). In the function space, differential operators and multiplication operators act locally on configuration spaces. From this observation, in addition to introducing $\Lambda(A),$ it may be suitable to assume some locality conditions in $U$. Therefore, in this paper, we impose “finite propagation” condition (see Assumption 2.3) for $U$. By these two ideas, we can analyze the commutator in detail.
The optimal constant $\delta$ in (1.1) depends on dispersion relations of quantum walks. For example, in [22, 25], the optimal constant is derived. However, in quantum walks, we can select graphs, internal degrees of freedom, motion of a quantum walker, and shift parameters. Thus, it would be useful to establish EDP in general settings. For example, in [37], Tiedra de Aldecoa considered spectral and scattering theory for quantum walks on not square lattices but trees. If discrete spectra of such quantum walks are not empty, we can apply our results. Our idea can be applied to discrete Schrödinger operators since they consist of shift operators and multiplication operators that act locally.
@@ 2. Set up and main result
  1. (1)
    $\sigma_{\mathrm{d}}(U)\neq\emptyset$.
  2. (2)
    The unitary operator $U$ finitely propagates with a constant $b>0$ with respect to a non-negative, possibly unbounded, self-adjoint operator $A.$
@@
Under Assunption 2.3, for any $\psi\in\mathrm{Ker}(U-\lambda)\setminus\{0\}$ with $\lambda\in\sigma_{\mathrm{d}}(U)$, $e^{\delta A}\psi\in\mathcal{H}$ for any $\delta>0$ such that $2\sinh(\delta b)<d(\lambda)$.
@@
The non-negativity in the second part of Assumption 2.3 is not essential. However, for simplicity, we assume the non-negativity of $A$ in this paper.
@@
$\|Uf-\lambda f\|\geq\{d(\lambda)-\epsilon\}\|f\|,$
@@
$a\|g_{R}\|$ $<\|Ug_{R}-\lambda g_{R}\|$
$\leq\|Uf_{R}-\lambda f_{R}\|+\|(U-\lambda)E_{U}(X)f_{R}\|$
$<d(\lambda)-\epsilon+2\|E_{U}(X)f_{R}\|.$
@@
$\Lambda(r):=\displaystyle\sum_{n=1}^{\infty}\delta nb\mathbb{I}_{B_{n}}(r),\quad\Lambda_{N}(r):=\begin{cases}\displaystyle\sum_{n=1}^{N}\delta nb\mathbb{I}_{B_{n}}(r),\ &r\in[0,Nb),\\ \delta Nb,&r\in[Nb,\infty),\end{cases}$ (2.1)
where $B_{n}:=[(n-1)b,nb)\subset\mathbb{R}$ and $\mathbb{I}_{B_{n}}$ is the characteristic function of $B_{n}.$ Then, $\Lambda$ approximates a function $f(r):=\delta r,\ (r\in[0,\infty))$ from the above and $\Lambda_{N}$ is a cut-off function of $\Lambda.$
@@
For any $R>0$, we set $E_{A}(R):=E_{A}([R,\infty)).$ Then, $e^{\Lambda(A)}[U,E_{A}(R)]$ is bounded on $\mathcal{H}$ and
$\|e^{\Lambda(A)}[U,E_{A}(R)]\|\leq e^{\delta\lceil R+b\rceil_{b}}+e^{\delta\lceil R\rceil_{b}},$
where for $x>0,\lceil x\rceil_{b}:=b\cdot\min\{n\in\mathbb{N}|\ x\leq nb\}.$
@@
where if $R-b\leq 0,$ we set $E_{A}([0,R-b))=0$ and $E_{A}([R-b,R))=E_{A}([0,R)).$ Thus, for any $\psi\in\mathcal{H},$ it follows that $[U,E_{A}(R)]\psi\in D(e^{\Lambda(A)})$ and
$\|e^{\Lambda(A)}[U,E_{A}(R)]\psi\|\leq\left(e^{\delta\lceil R+b\rceil_{b}}+e^{\delta\lceil R\rceil_{b}}\right)\|\psi\|.$
@@
$\|[U,e^{\Lambda_{N}(A)}]e^{-\Lambda_{N}(A)}\|\leq 2\sinh(\delta b).$
@@
The integrand in (2.2) is decomposed as follows:
$e^{t\Lambda_{N}(A)}[U,\Lambda_{N}(A)]e^{-t\Lambda_{N}(A)}$
$=e^{t\Lambda_{N}(A)}\{U\Lambda_{N}(A)-\Lambda_{N}(A)U\}E_{A}(B_{1})$
$+\displaystyle\sum_{m=2}^{N}e^{t\Lambda_{N}(A)}\{U\Lambda_{N}(A)-\Lambda_{N}(A)U\}e^{-t\Lambda_{N}(A)}E_{A}(B_{m})$
@@
$+e^{t\Lambda_{N}(A)}\{U\Lambda_{N}(A)-\Lambda_{N}(A)U\}e^{-t\Lambda_{N}(A)}E_{A}((N+1)b)$
$=:\mathrm{I}+\mathrm{II}+\mathrm{III}+\mathrm{IV}.$
@@
$=E_{A}(B_{2})e^{2t\delta b}(U\delta b-2\delta bU)e^{-t\delta b}E_{A}(B_{1})$
$=-\delta be^{t\delta b}E_{A}(B_{2})UE_{A}(B_{1}).$
@@
$\mathrm{II}$ $=\displaystyle\sum_{m=2}^{N}\{E_{A}(B_{m-1})+E_{A}(B_{m})+E_{A}(B_{m+1})\}$
×$\quad\times e^{t\Lambda_{N}(A)}\{U\Lambda_{N}(A)-\Lambda_{N}(A)U\}e^{-\Lambda_{N}(A)}E_{A}(B_{m})$
$=\displaystyle\sum_{m=2}^{N}\Big{[}E_{A}(B_{m-1})e^{t\delta b(m-1)}\{U\delta bm-\delta b(m-1)U\}e^{-t\delta bm}E_{A}(B_{m})$
$\quad+E_{A}(B_{m+1})e^{t\delta b(m+1)}\{U\delta bm-\delta b(m+1)U\}e^{-t\delta bm}E_{A}(B_{m})\Big{]}$
$=\delta b\displaystyle\sum_{m=2}^{N}\Big{[}e^{-t\delta b}E_{A}(B_{m-1})UE_{A}(B_{m})-e^{t\delta b}E_{A}(B_{m+1})UE_{A}(B_{m})\Big{]}.$
@@
$\mathrm{III}$ =$=\{E_{A}(B_{N})+E_{A}(B_{N+1})+E_{A}(B_{N+2})\}$
×$\quad\times e^{t\Lambda_{N}(A)}\{U\Lambda_{N}(A)-\Lambda_{N}(A)U\}e^{-t\Lambda_{N}(A)}E_{A}(B_{N+1})$
$=E_{A}(B_{N})e^{t\delta bN}\{U\delta b(N+1)-\delta bNU\}e^{-t\delta b(N+1)}E_{A}(B_{N+1})$
$=\delta be^{-t\delta b}E_{A}(B_{N})UE_{A}(B_{N+1}).$
Lastly, the forth term $\mathrm{IV}$ can be calculated as follows:
$\mathrm{IV}$ $=E_{A}(Nb)e^{t\Lambda_{N}(A)}\{U\Lambda_{N}(A)-\Lambda_{N}(A)U\}e^{-t\Lambda_{N}(A)}E_{A}((N+1)b)$
$=E_{A}(Nb)e^{t\Lambda_{N}(A)}(Ub\delta N-b\delta NU)e^{-t\Lambda_{N}(A)}E_{A}((N+1)b)$
@@
$[U,e^{\Lambda_{N}(A)}]e^{-\Lambda_{N}(A)}$
$=\delta b\displaystyle\int_{0}^{1}e^{-t\delta b}\mathrm{d}t\cdot\displaystyle\sum_{m=2}^{N+1}E_{A}(B_{m-1})UE_{A}(B_{m})-\delta b\displaystyle\int_{0}^{1}e^{t\delta b}\mathrm{d}t\cdot\displaystyle\sum_{m=1}^{N}E_{A}(B_{m+1})UE_{A}(B_{m})$
$=(1-e^{-\delta b})\displaystyle\sum_{m=2}^{N+1}E_{A}(B_{m-1})UE_{A}(B_{m})-(e^{\delta b}-1)\displaystyle\sum_{m=1}^{N}E_{A}(B_{m+1})UE_{A}(B_{m}).$
@@
$\|[U,e^{\Lambda_{N}(A)}]e^{-\Lambda_{N}(A)}\psi\|^{2}$
$=\|(1-e^{-\delta b})\displaystyle\sum_{m=2}^{N+1}E_{A}(B_{m-1})UE_{A}(B_{m})\psi-(e^{\delta b}-1)\displaystyle\sum_{m=1}^{N}E_{A}(B_{m+1})UE_{A}(B_{m})\psi\|^{2}$
$=(1-e^{-\delta b})^{2}\displaystyle\sum_{m=2}^{N+1}\|E_{A}(B_{m-1})UE_{A}(B_{m})\psi\|^{2}+(e^{\delta b}-1)^{2}\displaystyle\sum_{m=1}^{N}\|E_{A}(B_{m+1})UE_{A}(B_{m})\psi\|^{2}$
$-2(1-e^{-\delta b})(e^{\delta b}-1)\displaystyle\sum_{m=2}^{N+1}\displaystyle\sum_{n=1}^{N}\mathrm{Re}\langle E_{A}(B_{m-1})UE_{A}(B_{m})\psi,E_{A}(B_{n+1})UE_{A}(B_{n})\psi\rangle$
$=(1-e^{-\delta b})^{2}\displaystyle\sum_{m=2}^{N+1}\|E_{A}(B_{m-1})UE_{A}(B_{m})\psi\|^{2}+(e^{\delta b}-1)^{2}\displaystyle\sum_{m=1}^{N}\|E_{A}(B_{m+1})UE_{A}(B_{m})\psi\|^{2}$
$-2(1-e^{-\delta b})(e^{\delta b}-1)\displaystyle\sum_{n=2}^{N}\mathrm{Re}\langle E_{A}(B_{n})UE_{A}(B_{n+1})\psi,E_{A}(B_{n})UE_{A}(B_{n-1})\psi\rangle$
$\leq(1-e^{-\delta b})^{2}\displaystyle\sum_{m=2}^{N}\|E_{A}(B_{m})\psi\|^{2}+(e^{\delta b}-1)^{2}\displaystyle\sum_{m=1}^{N}\|E_{A}(B_{m})\psi\|^{2}$
$\quad+(1-e^{-\delta b})(e^{\delta b}-1)\displaystyle\sum_{n=2}^{N}\{\|E_{A}(B_{n})UE_{A}(B_{n+1})\psi\|^{2}+\|E_{A}(B_{n})UE_{A}(B_{n-1})\psi\|^{2}\}$
$\leq(1-e^{\delta b})^{2}\|\psi\|^{2}+(e^{\delta b}-1)^{2}\|\psi\|^{2}+2(1-e^{-\delta b})(e^{\delta b}-1)\|\psi\|^{2}$
=$=\{(1-e^{-\delta b})+(e^{\delta b}-1)\}^{2}\|\psi\|^{2}$
$=(e^{\delta b}-e^{-\delta b})^{2}\|\psi\|^{2}.$
@@
We choose $\epsilon>0$ as $\epsilon:=[d(\lambda)-2\sinh(\delta b)]/2.$ Then, by Lemma 2.6, there exists $R>0$ such that for any $f\in\mathcal{H}$ with $E_{A}([0,R))f=0,$ we have
@@
We take $\psi\in\mathrm{Ker}(U-\lambda)\setminus\{0\}$ with $\lambda\in\sigma_{\mathrm{d}}(U)$. For $R$ and $b$, there exists $N_{0}\in\mathbb{N}$ such that $R<N_{0}b$. Then we set $f_{N}:=e^{\Lambda_{N}(A)}E_{A}(R)\psi,\ (N\geq N_{0}).$ Since $E_{A}([0,R))f_{N}=0$, we have the following for arbitrary $N\geq N_{0}$:
@@
$Uf_{N}-\lambda f_{N}=[U,e^{\Lambda_{N}(A)}E_{A}(R)]\psi=[U,e^{\Lambda_{N}(A)}]E_{A}(R)\psi+e^{\Lambda_{N}(A)}[U,E_{A}(R)]\psi.$ (2.4)
@@
$\|e^{\Lambda_{N}(A)}[U,E_{A}(R)]\psi\|\leq\left(e^{\delta\lceil R+b\rceil_{b}}+e^{\delta\lceil R\rceil_{b}}\right)\|\psi\|,$
For the first term of (2.4), from Lemma 2.8, we get
$\|[U,e^{\Lambda_{N}(A)}]E_{A}(R)\psi\|=\|[U,e^{\Lambda_{N}(A)}]e^{-\Lambda_{N}(A)}e^{\Lambda_{N}(A)}E_{A}(R)\psi\|\leq 2\sinh(\delta b)\|f_{N}\|.$
@@
$\|Uf_{N}-\lambda f_{N}\|\leq\left(e^{\delta\lceil R+b\rceil_{b}}+e^{\delta\lceil R\rceil_{b}}\right)\|\psi\|+2\sinh(\delta b)\|f_{N}\|.$
From the above inequality and (2.3), we arrive at
$\displaystyle\frac{d(\lambda)-2\sinh(\delta b)}{2}\|f_{N}\|\leq\left(e^{\delta\lceil R+b\rceil_{b}}+e^{\delta\lceil R\rceil_{b}}\right)\|\psi\|.$ (2.5)
Since $N$ is arbitrary and right hand side of (2.5) is independent of $N$, we conclude that $e^{\Lambda(A)}\psi\in\mathcal{H}$ by the monotone convergence theorem. This implies $e^{\delta A}\psi\in\mathcal{H}.$
3. Application